# Causality-enhanced Decision-Making for Autonomous Mobile Robots in Dynamic Environments

### v3

@@ 5.5 Discussions
Fig. 9 (top-left) shows that TIAGo, employing our causality-based decision-making framework, successfully completes $89\%$ of the tasks (1860 out of 2090), compared to only $55.9\%$ for the baseline. This difference is also reflected in the failure rates due to congestion (orange) and battery depletion (red). Specifically, congestion-related failures account for $10.9\%$ in our approach versus $43.8\%$ in the baseline. This demonstrates that the estimations of people density ($D$) and battery consumption ($L$) enable the robot to complete nearly $90\%$ of the tasks, whereas the baseline, lacking this knowledge, fails in almost half of the cases. An example of the experiments in Fig. 11 shows a typical scenarios where TIAGo, using our approach, successfully avoids congested areas and completes its task. In contrast, the baseline approach, which prioritises the shortest path—even when it is the busiest—leads the robot into congestion, ultimately causing task failure.
Although the effect is smaller, our approach also improves task-execution decisions based on estimated battery consumption. The causal approach successfully estimates in 7 out of 10 cases whether the battery level is sufficient to complete a task without reaching the critical threshold. In contrast, the baseline approach succeeds only 3 times out of 10, as the robot executes tasks regardless of its battery level and task cost.
Figure 10: Summary of overall safety results across all time slots. (Left) Number of human-robot collisions: (blue) baseline approach (orange) our approach. (Right) Human-robot proxemic compliance with proxemic zone bands: (green) public, (blue) social, (orange) personal, (red) intimate.
@@
Figure 11: (Top-left) TIAGo robot with a person in a real-world scenario. (Top-right) TIAGo robot with a simulated agent in PeopleFlow. (Centre and bottom rows) Visualisations of the TIAGo robot executing its assigned tasks in two scenarios: (left) $S2$ working time and (right) $S6$ lunchtime. The robot’s goal is represented by a green sphere, while the planned path is highlighted in blue. Congested areas align with Fig. 7 in both scenarios. (centre) Path generated by the shortest-strategy heuristic in the baseline approach. (bottom) Path generated by our approach. The bottom-line graphs illustrate the effectiveness of our method, which prioritises safety over minimising travel distance.
@@
The overall distance travelled by the robot is illustrated in Fig. 9 (top-right). The impact of these path selection strategies is evident in the wasted distance (red blocks), which represents the distance travelled in tasks that ultimately failed. The baseline approach, which relies on the shortest-path strategy, results in $46.5\%$ of the total travelled distance being wasted. In contrast, our causality-based approach reduces this to only $10.2\%$, demonstrating a significantly more efficient motion. The sum of the planned distance (blue) and the extra distance (orange) represents the “effective” distance travelled to complete successful tasks. While the total travelled distance is slightly higher in the causal approach, it enables the robot to use about $90\%$ of the total travelled distance for successful tasks, compared to only $53.6\%$ for the baseline case.
These findings are consistent with the task time results in Fig. 9 (bottom-left). Specifically, the robot actively moves toward its goals for $84.9\%$ of the total working hours using our approach, compared to only $54.3\%$ for the baseline. Moreover, the wasted time—the time spent on tasks that ultimately failed—is drastically reduced from $38\%$ in the baseline to just $6.7\%$ in the causal case.
Finally, the battery usage in Fig. 9 (bottom-right) shows that the causal approach requires an almost equivalent number of battery cycles ($12$ for causal and $11$ for baseline), although many more tasks were successfully completed by our approach. Indeed, the battery usage is significantly better managed by the latter, with $92.3\%$ of the battery consumption classified as effective—i.e., spent on tasks that ultimately succeed—while only $7.7\%$ is wasted on failed tasks. In contrast, the baseline approach wasted $39.9\%$ of its battery on failed tasks, which corresponds to more than 4 battery cycles.
@@
Note that, in $S11$ (off-time), both methods behave identically. Specifically, in both the baseline and causal approaches, the robot successfully completes all 90 tasks, travelling comparable distances in the same amount of time and using the same amount of battery. This demonstrates that our approach does not introduce unnecessary path deviations in the absence of workers and congestion.
@@
Fig. 10 presents the safety-related results. The bar chart (left), which shows the number of times the robot and humans collides, reveals a significant difference between the two approaches. The baseline (blue bar), without information about $D$ and $L$, relies solely on the shortest-path heuristic. This limitation forces the robot to select the shortest but often busiest path (as shown in Fig. 11), leading to 182 collisions. In contrast, our approach (orange bar), which takes into account people density and battery consumption, leads to only 27 collisions in total.
This observation is further confirmed by the overall distribution of robot-human distances in Fig. 10 (right). This also highlights how both the baseline and causal approaches comply with proxemic zones [68]. In both approaches, the median distance falls within the public zone (green), indicating that on average, the robot maintains a safe distance from humans. However, a key difference emerges when examining the interquartile range (IQR) and lower quantiles. In the baseline approach, the 25th–75th percentile range extends into the social zone (blue), meaning that the robot frequently operates at closer distances to people. More critically, the lower whisker shows that the robot often enters the personal (orange) and intimate (red) zones, indicating also frequent close interactions. Our approach instead maintains the entire IQR within the public zone. The lower whisker and outliers rarely reach the personal and intimate zones, suggesting that close encounters are significantly reduced compared to the baseline. This further highlights the effectiveness of our causality-based decision framework in ensuring that the robot operates at safer distances.
@@
6 Causal vs Bayesian Inference
Although the advantages of causal inference over a standard Bayesian approach are well established in the literature [48], in this section we demonstrate such benefits through a simplified ablation study focused on our mobile robot’s decision-making process.
Figure 12: Simplified scenario. (Top) The TIAGo robot navigating between its target positions to complete tasks while avoiding unexpected obstacles (red cubes) in the environment. (Bottom-left) Causal model retrieved from the preliminary study. Context variables are shown in grey, while system variables are highlighted in orange. The thickness of the arrows and node borders represents the strength of cross- and auto-causal dependencies, respectively, while the numbers on each node and link indicate the time lag of the dependency. (Bottom-right) Causal inference engine derived from the built causal model.
Simplified Scenario
In this study we focus only on the estimation of $L$ using the causal inference engine. To simplify the analysis in terms number of variables involved, we do not consider people congestion areas. In this simplified scenario the TIAGo robot navigates between its target positions (diamond symbols in Fig. 4, top) while draining the battery according to the model in (1) in the discharging case ($C=0$). Additionally, we introduce unexpected obstacles (e.g., red box in Fig. 12 top) that have an impact on robot velocity and battery consumption.
The robot continuously moves between target stations, simulating the delivery of items from trays to conveyors. During task execution, random obstacles appear in its path, with a pre-defined probability. These force the robot to adjust its velocity to avoid collisions and overtake them to complete its task. Once avoided, the obstacle disappears.
We simulate 1 hour of operation, extracting from our simulator $V$, $L$, and the binary contextual factor $O$. We assumed the robot starts with a fully charged battery (${B=100}$) and set the battery parameters by aligning their values with the actual battery lifetime7 in both idle and motion modes. The recorded time-series is then subsampled and stored by the data pipeline, preparing it for the following learning pipeline.
The causal model, illustrated in Fig. 12 (bottom-left), is discovered by the J-PCMCI+ algorithm, chosen because of the presence of the context variable $O$, assuming a first-order Markov process (one time step lag). As expected, the model is essentially a simplified version of the previous one in Fig. 5, restricted to this specific case where only the $V$, $L$, and $O$ variables are considered ($L_{t}~{}\xleftarrow{}~{}C_{t}~{}\xrightarrow{}~{}V_{t}$ is missing since we consider only $C=0$). The final discretised model resulting from the parameter learning step is shown in Fig. 12 (bottom-right).
Theoretical Analysis
When estimating the relationship between two variables in the presence of a confounder, standard Bayesian inference operates purely in an observational setting, modelling the joint probability distribution and updating beliefs based on conditional dependencies. However, Bayesian networks do not distinguish between correlation and causation, making them ineffective when trying to determine the causal impact of one variable on another. This limitation becomes critical when a confounder influences both the treatment and the outcome, introducing spurious associations that bias the estimation. In contrast, causal inference explicitly models interventions using the do-operator, which allows us to estimate the cause-effect while accounting for confounders.
As explained in Sec. 4.1, the factor $O$ represents unexpected static obstacles in the environment, which influence both the robot’s velocity $V$ and battery consumption $L$. This makes $O$ a confounder between these two variables. In our scenario, as outlined in Sec. 3.3, the robot is assigned a task and queries the causal inference engine to make a decision whether to start or abort the task based on its battery level. Hence, an intervention on $V$ is performed to estimate $L$ before the task begins. However, the robot cannot know in advance whether it will encounter an unexpected obstacle along its path. Thus, we cannot directly condition on the confounder $O$, since its value is unknown at the time of decision-making. As a result, failing to account for this dependency would bias the causal effect estimation. Instead, by adjusting for $O$, we can isolate the true causal relationship between $V$ and $L$, ensuring a more accurate estimate of battery consumption in an uncertain scenario.
A typical Bayesian approach would estimate the probability of battery consumption at time $t$, given a certain robot velocity, applying the law of total probability with conditioning as follows:
$\text{P}(L_{t}\mid V_{t})=\sum_{O_{t}}\text{P}(L_{t}\mid V_{t},O_{t})\text{P}(O_{t}\mid V_{t})$ (11)
This formulation does not perform any intervention, i.e., it selects a particular $V$ without breaking all the incoming edges. By conditioning on the observed velocity, ${\text{P}(O_{t}\mid V_{t})}$ implicitly selects specific values of $O$ depending on $V$. Such indirect selection introduces a confounding bias, since the estimated effect of $V$ on $L$ remains influenced by $O$.
To correctly estimate the causal effect of $V$ on $L$, we must apply the do-operator to simulate an intervention:
$\text{P}(L_{t}\mid do(V_{t}))=\sum_{O_{t}}\text{P}(L_{t}\mid V_{t},O_{t})\text{P}(O_{t})$ (12)
The equation is essentially a backdoor adjustment [48], where we account for the confounder $O$ by integrating over its distribution. This removes the confounding bias, yielding an unbiased causal effect of $V$ on $L$.
Numerical Analysis
By simulating numerically a real query, we can better appreciate the difference between the two methods. Consider the following scenario inspired by a typical task assigned to the robot during our experiments presented in Sec. 5: TIAGo is at point A and must deliver an object to point B, with a distance of 60m between them. There is a 30% probability of some unexpected obstacle $O$ obstructing its path during the task. The robot’s query is “what if I cover this distance at my maximum velocity?”. At its maximum velocity—0.5m/s for the TIAGo robot—TIAGo will need ${\Delta_{t}=120}$ seconds to cover the distance, used in (1) to calculate the reference value of $L$. Since the value of $O$ is unknown at decision time, the correct reference for evaluation is the expected value of $L$, averaging over the two possible outcomes (i.e., ${O=0}$ and ${O=1}$). This yields a reference battery consumption of $L_{\text{REF}}=-2.504$, where REF stands for reference. Using Eq. (11) for Bayesian inference and Eq. (12) for causal inference, we estimate the probability distributions of $L$ under conditioning (Bayesian setting) and intervention (causal setting), respectively. By computing the expectation, we obtain the numerical estimates for the Bayesian network $\hat{L}_{\text{BN}}=-1.486$ and causal model $\hat{L}_{\text{CM}}=-2.142$, with the corresponding absolute errors from the reference value are ${\epsilon_{BN}=~{}\mid L_{\text{REF}}-\hat{L}_{\text{BN}}\mid~{}=1.018}$ and ${\epsilon_{CM}=~{}\mid L_{\text{REF}}-\hat{L}_{\text{CM}}\mid~{}=0.362}$. This shows that the causal model estimates the battery consumption with an error nearly three times smaller than the Bayesian network.
Figure 13: Comparison of the $\hat{L}$ estimation error for the query “What if the mobile robot covers a distance of 60m at velocity $v$?”. The bars represent the MAE across different $O$ probabilities, ranging from 10% to 50%, for the two inference methods: (blue) Bayesian inference and (orange) Causal inference. Error bars indicate the standard deviation of the absolute error.
To generalise, we varied the obstacle probability between $10\%$ and $50\%$, using $10\%$ increments, and compared the two inference methods. Fig. 13 shows the mean absolute error (MAE), for the two inference methods—(blue) Bayesian inference and (orange) Causal inference— across different $O$ probabilities (i.e., $10\%$, $20\%$, $30\%$, $40\%$, $50\%$) with error bars representing standard deviations. This comparison demonstrates that, on average, the causal model outperforms the Bayesian network in inference accuracy across the various obstacle probabilities.
@@ 7 Conclusion
In this work, we introduced a novel causality-based decision-making framework and a new HRSI simulator featuring a mobile robot, multiple pedestrians, and various contextual factors. These factors influence the dynamics of human and robot’s goals, their interactions, and the robot’s task execution. Our causality framework integrates data processing, causal discovery, and causal inference, enabling the robot to make better path planning decisions and to determine in advance whether to proceed with or cancel a task. These decisions are based on two causally-inferred quantities: expected people densities in specific areas of the environment, and estimated battery consumption to reach a target location. The proposed system was applied and evaluated in a simulated warehouse scenario with context-dependent goals and interactions, made with PeopleFlow. The results validate the improvements achieved by our solution in terms of planning efficiency and safety for shared human-robot environments.
Building on these results, future work will focus on deploying our approach on a real robot for onboard online decision-making. This step is crucial for evaluating the effectiveness of our method in real-world logistics scenarios involving multiple humans, shared spaces, and complex interactions. Additionally, a promising direction is to enhance our causality-based decision-making framework by integrating a module for causal feature learning (i.e., which variables are important for predicting a certain outcome) into the learning pipeline. Beyond learning the causal structure and parameters, this addition would enable the robot to automatically identify and refine relevant features, enhancing generalisation and adaptability in dynamic environments.
Acknowledgments

### v4

@@ 5.5 Discussions
Fig. 10(a) now clearly illustrates the benefits of our framework and the distinct contributions of its components. The Full Causal framework achieves the highest success rate at $89.0\%$ ($1860$ out of $2090$ tasks), a substantial improvement compared to the baseline’s $55.9\%$. The ablation study shows the source of this gain is due to the use of exploiting the causal inference engine for the path planning (Eq. 9). Indeed, the Causal Routing approach is clearly the primary driver of improved performance. By itself, it boosts the success rate to $87.7\%$, almost entirely by reducing congestion failures (Failures D) from $43.8\%$ (Baseline) to $11.8\%$. In contrast, the Refusal-Only approach, which still uses the shortest-path router, provides almost no benefit ($57.5\%$ success), proving that simply refusing tasks is not an effective strategy when the routing itself is poor. However, Causal Routing is still vulnerable to battery-related failures, which occur in $0.5\%$ of the tasks, while the Refusal-Only approach, which has the refusal mechanism, suffers only $0.2\%$ battery failures. This shows a clear synergistic benefit in our Full Causal framework: by combining smart routing with the refusal mechanism, it catches the edge cases, reducing battery failures to just $0.1\%$ and achieving the highest overall success rate. This demonstrates that both estimations of people density ($D$) and battery consumption ($L$) are necessary, with the causal heuristic eliminating the majority of congestion failures and the refusal mechanism providing a critical safeguard against battery-related failures. An example of the experiments in Fig. 12 shows a typical scenarios where TIAGo, using our approach, successfully avoids congested areas and completes its task. In contrast, the baseline approach, which prioritises the shortest path—even when it is the busiest—leads the robot into congestion, ultimately causing task failure.
@@
Figure 12: (a) TIAGo robot with a person in a real-world scenario. (b) TIAGo robot with a simulated agent in PeopleFlow. (c, d, e and f) Visualisations of the TIAGo robot executing its assigned tasks in two scenarios: (c, e) $S2$ working time and (d, f) $S6$ lunchtime. The robot’s goal is represented by a green sphere, while the planned path is highlighted in blue. Congested areas align with Fig. 7 in both scenarios. (c, d) Path generated by the shortest-strategy heuristic in the baseline approach. (e, f) Path generated by our approach. The bottom-line graphs illustrate the effectiveness of our method, which prioritises safety over minimising travel distance.
@@
The overall distance travelled by the robot is illustrated in Fig. 10(b). impact of these path selection strategies is evident in the wasted distance (red blocks), which represents the distance travelled in tasks that ultimately failed. The Baseline approach, which relies on the shortest-path strategy, results in $46.5\%$ of its total travelled distance being wasted. The ablation study clearly shows that the Causal Routing heuristic is the primary solution to this: by itself, it drastically reduces wasted distance to just $11.1\%$. In contrast, the Refusal-Only approach still wastes $37.4\%$ of its distance, once again proving that smart refusal is ineffective without smart routing. The Full Causal framework combines both, achieving the lowest wasted distance of all at $10.2\%$, demonstrating a significant improvement in motion efficiency.
These findings are consistent with the task time results in Fig. 10(c). Specifically, the wasted time (red blocks)—the time spent on tasks that ultimately failed—is greatly reduced. The Baseline approach wastes $37.7\%$ of its total time on failed tasks. The ablation study again shows the Causal Routing heuristic is the primary driver of performance, by itself drastically reducing wasted time to just $6.6\%$. The Refusal-Only approach, in contrast, shows almost no improvement, remaining at $35.9\%$ wasted time. The Full Causal framework ($6.6\%$ wasted) performs identically to the Causal Routing approach, confirming that the smart heuristic is the main source of this efficiency gain. This reduction in waste directly translates to a massive increase in productive work. The robot actively moves toward its goals (blue block) for $83.8\%$ and $84.6\%$ of the total time in the Full Causal framework and in Causal Routing, respectively, compared to only $53.6\%$ for the Baseline and $56\%$ for the Refusal-Only.
Finally, the battery usage in Fig. 10(d) highlights the framework’s impact on energy efficiency. The Baseline approach wastes a significant $39.9\%$ of its total battery consumption on failed tasks. The ablation study clearly shows that the Causal Routing heuristic is the primary solution to this problem. By itself, it reduces wasted battery to just $8\%$. In contrast, the Refusal-Only approach is aligned with the baseline result, wasting $37.0\%$ of its battery. The Full Causal framework achieves the best result, wasting only $7.7\%$ of its battery. This slight difference over Causal Routing ($8.0\%$) is important: it confirms the synergistic advantage of the refusal mechanism. As shown in Fig. 10(a), the refusal component catches the few catastrophic battery failures (Failures L) that Causal Routing cannot preemptively avoid, making the Full Causal framework the most robust. This is a key finding: our framework’s battery usage is significantly better managed, dedicating $92.3\%$ of its consumption to successful tasks, compared to only $60.1\%$ for the baseline.
@@
Note that, in $S11$ (off-time), both methods behave identically. Specifically, in all approaches, the robot successfully completes all $90$ tasks, travelling comparable distances in the same amount of time and using the same amount of battery. This demonstrates that our approach does not introduce unnecessary path deviations in the absence of workers and congestion.
@@
Beyond task-level efficiency, we also measured the computational cost of the inference pipeline. The average inference time per task was approximately $0.32$ s, as detailed in Table 1. The table compares two runtime performance metrics: the average query time and the average number of candidate routes per task across all time slots. In the causal case, the number of candidate paths evaluated by the heuristic function is slightly higher compared to the baseline approach. The query time for the baseline is, by definition, 0 s, whereas for the causal approach it is $0.327$ s. This confirms that the computational overhead introduced by our causal query remains low enough for practical, real-time planning. It is important to note that this time was recorded on hardware that was simultaneously running the inference code and the computationally expensive Gazebo physics simulation. On a real robot’s onboard computer, which would only run the inference, this query time would be lower, confirming its suitability for real-time planning. As demonstrated in our scalability analysis (see A), the query cost scales linearly with the size of the map. This predictable, linear behaviour confirms its suitability for real-time planning applications, as the cost does not explode with map complexity.
@@
Fig. 11 presents the safety-related results, where the ablation study highlights the source of our framework’s benefits. Fig. 11(a) reveals a substantial improvement in the human-robot collision metric. The Baseline approach, relying on the shortest-path, is the most dangerous, resulting in $194$ collisions. The ablation shows that the Refusal-Only approach performs almost identically, with $187$ collisions. This proves that simply refusing tasks does not improve safety. The safety gain comes entirely from the Causal Routing heuristic: by itself, it reduces collisions to just $80$. The Full Causal framework is marginally better at $78$ collisions, confirming that the smart heuristic—which proactively routes the robot away from busy areas—is the primary driver of safety.
This observation is confirmed by the proxemic compliance in Fig. 11(b). The Baseline and Refusal-Only approaches are again nearly identical: their interquartile ranges (IQRs) dip into the social zone, and their lower whiskers show frequent, close interactions in the personal and intimate zones. In contrast, both the Causal Routing and Full Causal approaches are far safer. Their median distances are higher, and their entire IQRs are maintained within the public zone. This demonstrates that the causal heuristic (Eq. 9) is highly effective in ensuring the robot operates at safer distances.
@@
To better demonstrate the tangible benefits of our causal approach, a supplementary video showing the experimental environment created by PeopleFlow and comparing the baseline and full causal approaches in two scenarios is available online [^15] https://drive.google.com/file/d/1oBtTpPZbstD2a66BGxG7XdfBZ84MLPM7/view?usp=sharing.
@@ 6 Conclusion
In this work, we presented a novel end-to-end, causality-based decision-making framework that, for the first time, combines causal discovery from sensor data with causal reasoning for motion planning in ROS. We also introduced PeopleFlow, a new HRSI simulator featuring a mobile robot, multiple pedestrians, and diverse contextual factors that influence human and robot goals, their interactions, and the robot’s task execution. Our approach unifies data processing, causal discovery, and causal inference into a single pipeline, enabling the robot to make better path planning decisions and to determine in advance whether to proceed with or cancel a task. These decisions are based on two causally-inferred quantities: expected people densities in specific areas of the environment, and estimated battery consumption to reach a target location. We evaluated the proposed system in a simulated warehouse scenario, created using PeopleFlow. The scenario, inspired by our funding project DARKO, incorporates both workers and a robot with context-dependent goals, where the spatial behaviour of each is mutually influenced by the presence of the other. The results validate the improvements achieved by our solution in terms of planning efficiency and safety for shared human-robot environments.
Building on these results, our primary future work will focus on real-world validation. This involves deploying the framework on a real robot for onboard online decision-making, a crucial step for evaluating its effectiveness in logistics scenarios involving multiple humans and complex interactions. We will also explore the key advantage of our end-to-end architecture: its potential for long-term autonomy. This includes enabling the robot to autonomously re-run the learning pipeline on new data to adapt to environmental changes, such as new worker schedules, without human intervention. Another important direction is multi-robot scalability. While our current model is single-robot, the framework is designed to scale by allowing a central fleet manager to use our causal model as a component in a multi-agent heuristic for conflict-resolution. Finally, a promising direction is to enhance the learning pipeline by integrating a module for causal feature learning (i.e., which variables are important for predicting a certain outcome), which would enable the robot to automatically identify and refine relevant features, further enhancing its adaptability in dynamic environments.
Acknowledgements
@@
Appendix A Causal Query Scalability Analysis
Figure 13: Causal query scalability analysis.
A critical requirement for any online planning system is a predictable and efficient runtime. A complex query for choosing a path, even if accurate, becomes impractical if it introduces multi-second delays or has a non-linear (e.g., exponential) cost that makes it unusable on larger maps. Therefore, we conducted a detailed scalability analysis to rigorously evaluate the computational cost of our causal query and verify its suitability for real-time applications.
Beyond the hardware configuration (specified in Sec. 5), it is important to understand how the causal query scales with the number of waypoints. Therefore, we incrementally tested the system by generating sub-graphs of increasing size, starting from $10$ waypoints and increasing in steps of $10$ up to the full map ($73$ waypoints). For each sub-graph size, we executed the full prediction query $1000$ times, each with a randomised time slot and robot velocity.
The result of this scalability analysis is shown in Fig. 13, where we report the mean and standard deviation of the inference time for each sub-graph size. This analysis revealed a highly predictable linear relationship between the number of waypoints and the mean inference time. The error bars, representing the standard deviation, confirm the consistency of the query time. This predictable, linear scaling behaviour confirms that the computational cost does not explode with map complexity, which allows us to confidently predict performance even on larger maps and validates the approach for real-time planning.
Appendix B Heuristic Weight Sensitivity Analysis
To evaluating the impact of heuristic weight changes on decision outcomes, we performed a comprehensive sensitivity analysis. The objective was to systematically validate our parameter selection and confirm that it represents an optimal trade-off between task efficiency and robot safety.
B.1 Methodology
To conduct this analysis, we evaluated $5$ representative delivery tasks. These tasks were executed twice, during two distinct time slots (H2, a working time, and H6, a lunch time), to assess the framework’s performance in two different scenarios of human activity.
We selected three distinct levels for each of the three heuristic weights: $\lambda_{\delta}$ ($0.1$, 11, $10$), $\lambda_{D}$ (11, $10$, $100$), and $\lambda_{L}$ ($0.5$, 55, $50$). This resulted in $27$ unique weight configurations, allowing us to systematically study their interactions and impact on performance. This selection of logarithmically-spaced values was chosen to test a wide range of magnitudes to thoroughly validate our intended prioritisation strategy of people ($\lambda_{D}$) over battery cost ($\lambda_{L}$), and finally over goal distance ($\lambda_{\delta}$).
B.2 Results and Discussion
The results of this analysis are summarised in Table 2. As hypothesised, the analysis confirms that the configurations yielding the best performance, under the combined efficiency and safety metrics defined in Sec. 5.3, were indeed those that adhered to this intended prioritisation.
From Table 2, we first filtered these $27$ configurations to ensure robust and safe performance. Any configuration that resulted in one or more collisions (Collis. >0>0) or failed to achieve a $100\%$ mission success rate (Succ. $100\%$) was excluded from consideration (marked in red). From the remaining set of successful and safe configurations, we observed that configurations maintaining this $\lambda_{D}>\lambda_{L}>\lambda_{\delta}$ order consistently produced the strongest results. Although all configurations within this successful group yielded broadly similar, high-quality performance, we selected the parameters ($\lambda_{\delta}=1$, $\lambda_{D}=10$, $\lambda_{L}=5$) (marked in green) as the optimal set. This configuration, ($1$, $10$, 55), offers the best trade-off between operational performance (i.e., task time, travelled distance, and battery usage) and safety, maintaining a significantly larger minimum distance of $0.53$ m. Overall, this parameter combination represents the optimal balance between task efficiency and human safety.
Table 2: Comprehensive sensitivity analysis of $\lambda_{\delta}$, $\lambda_{D}$, and $\lambda_{L}$ Parameters
Configuration Succ. Collis. Time Distance Battery Proxemics
($\lambda_{\delta},\lambda_{D},\lambda_{L}$) Tot. Active Stalled Wasted Tot. Planned Extra Wasted Tot. Effective Wasted Median Q1 Q3 Min
(%) (count) (h) (%) (%) (%) (m) (%) (%) (%) (cycles) (%) (%) (m) (m) (m) (m)
(0.1, 1, 0.5) 100.0 0 0.198 96.2 3.8 0.0 0.269 93.6 6.4 0.0 0.045 100.0 0.0 5.80 4.70 6.76 0.39
(0.1, 1, 5) 70.0 0 0.198 63.5 4.0 32.4 0.231 67.1 4.9 28.0 0.044 67.2 32.8 5.04 3.62 6.17 0.32
(0.1, 1, 50) 40.0 5 0.272 46.0 8.9 45.1 0.371 23.5 10.4 66.1 0.057 50.3 49.7 4.48 2.97 5.89 0.07
(0.1, 10, 0.5) 100.0 0 0.197 96.4 3.6 0.0 0.264 95.4 4.6 0.0 0.045 100.0 0.0 5.78 4.72 6.76 0.37
(0.1, 10, 5) 100.0 0 0.209 95.0 5.0 0.0 0.278 90.3 9.7 0.0 0.047 100.0 0.0 5.79 4.71 6.76 0.37
(0.1, 10, 50) 70.0 0 0.199 64.7 4.0 31.4 0.234 66.2 8.6 25.3 0.044 68.4 31.6 4.81 3.44 6.09 0.37
(0.1, 100, 0.5) 100.0 0 0.203 97.0 3.0 0.0 0.280 96.5 3.5 0.0 0.047 100.0 0.0 5.55 4.49 6.58 0.31
(0.1, 100, 5) 100.0 0 0.204 96.2 3.8 0.0 0.276 96.0 4.0 0.0 0.047 100.0 0.0 5.63 4.61 6.64 0.58
(0.1, 100, 50) 100.0 1 0.217 95.8 4.2 0.0 0.297 89.1 10.9 0.0 0.050 100.0 0.0 5.67 4.69 6.67 0.22
(1, 1, 0.5) 100.0 1 0.215 94.6 5.4 0.0 0.289 87.0 13.0 0.0 0.048 100.0 0.0 5.72 4.60 6.68 0.26
(1, 1, 5) 70.0 3 0.183 67.3 3.2 29.5 0.223 69.5 4.4 26.1 0.041 70.2 29.8 5.07 3.62 6.23 0.09
(1, 1, 50) 50.0 2 0.243 51.0 9.0 40.0 0.295 37.6 8.8 53.5 0.050 56.5 43.5 4.47 3.01 5.97 0.20
(1, 10, 0.5) 100.0 0 0.191 96.8 3.2 0.0 0.261 96.2 3.8 0.0 0.044 100.0 0.0 5.78 4.73 6.77 0.47
(1, 10, 5) 100.0 0 0.193 96.8 3.2 0.0 0.262 96.0 4.0 0.0 0.044 100.0 0.0 5.75 4.71 6.78 0.53
(1, 10, 50) 80.0 1 0.201 75.8 4.7 19.4 0.251 70.1 12.1 17.8 0.045 80.0 20.0 4.98 3.65 6.14 0.27
(1, 100, 0.5) 100.0 1 0.206 95.2 4.8 0.0 0.275 91.5 8.5 0.0 0.047 100.0 0.0 5.76 4.69 6.78 0.08
(1, 100, 5) 100.0 0 0.192 97.2 2.8 0.0 0.267 94.2 5.8 0.0 0.045 100.0 0.0 5.83 4.75 6.78 0.52
(1, 100, 50) 100.0 0 0.199 95.7 4.3 0.0 0.267 94.4 5.6 0.0 0.045 100.0 0.0 5.85 4.75 6.86 0.42
(10, 1, 0.5) 40.0 0 0.229 44.1 8.7 47.1 0.198 40.4 15.8 43.8 0.047 49.0 51.0 4.38 2.93 5.88 0.35
(10, 1, 5) 40.0 2 0.203 41.3 2.7 56.1 0.214 39.0 9.0 52.0 0.045 43.2 56.8 4.36 2.93 5.91 0.12
(10, 1, 50) 10.0 8 0.188 8.8 0.3 90.9 2.864 0.6 0.1 99.3 0.041 9.3 90.7 4.56 3.02 5.91 0.01
(10, 10, 0.5) 100.0 0 0.202 95.1 4.9 0.0 0.263 94.4 5.6 0.0 0.046 100.0 0.0 5.92 4.85 6.85 0.32
(10, 10, 5) 100.0 0 0.189 97.4 2.6 0.0 0.262 95.9 4.1 0.0 0.044 100.0 0.0 5.76 4.65 6.76 0.45
(10, 10, 50) 60.0 0 0.196 52.7 1.0 46.3 0.222 60.1 4.6 35.3 0.045 54.8 45.2 4.87 3.38 6.06 0.46
(10, 100, 0.5) 100.0 0 0.199 96.6 3.4 0.0 0.271 91.9 8.1 0.0 0.046 100.0 0.0 5.84 4.86 6.77 0.29
(10, 100, 5) 90.0 1 0.186 88.1 3.0 8.9 0.253 87.6 3.5 8.8 0.043 91.1 8.9 5.77 4.74 6.78 0.10
(10, 100, 50) 100.0 0 0.192 96.6 3.4 0.0 0.262 96.1 3.9 0.0 0.044 100.0 0.0 5.73 4.70 6.82 0.51